A popular resort hotel has 600 rooms and is usually fully booked. About 5​% of the time a reservation is canceled before the​ 6:00 p.m. deadline with no penalty. What is the probability that at least 564 rooms will be​ occupied? Use the binomial distribution to find the exact value.
Answer #1

Solution:-

Given that :-

p = 0.05

n = 600

P( X \geq 564 )

= 1 - P (X \leq 564)

= 1 - BINOM.DIST( 564,600,0.05, TRUE)

=1- 1

=0

P(X \ \geq 564) =0